# Neurons Speak in Ranges: Breaking Free from Discrete Neuronal Attribution

### v1

@@ 4 Neuronal Activation Patterns
To quantitatively assess normality, while accounting for practical significance, we employ the KS test with an effect size threshold of 10%. This approach tests whether the distribution remains within a reasonable bound of normality, rather than testing for perfect normality, which is overly strict for real-world data. For each neuron, we normalize the activations to zero mean and unit variance, then compute the KS statistic against a standard normal distribution. The KS statistic represents the maximum absolute difference between the empirical and theoretical cumulative distribution functions. Using a threshold of 0.1 (allowing maximum 10% deviation from normal), we find that close to 100% of the neurons exhibit practically normal distributions. The combination of near-ideal skewness and kurtosis values, visual confirmation through KDEs, and our effect size-based KS tests provide strong evidence that the activations follow approximately normal distributions.
Table 3: Evaluation of selected models on IMDB, SST2, AG-News, and DBPedia-14 datasets using activation range and neuron masking techniques. Performance metrics, calculated using trimmed means at the class level, include class accuracy (Acc), class prediction probability (Conf), average accuracy across other classes (CAcc), and average class prediction probability across other classes (CConf). Base Values represent baseline model performance, Activation Range Masking and Neuron Masking entries show the drop from these baselines. The trimmed mean is calculated by removing extreme values (top and bottom 10%) before computing the average to reduce the impact of outliers. We only bold the drops in auxiliary concepts (CAcc, CConf). The Drops in targeted concepts are comparable with Neuron Masking with slightly better results overall. For Bert and GPT-2 50% and for Llama-3.2-3B 30% neurons selected.
@@
5 Activation Ranges-guided Intervention
Given that the neuronal activations are Gaussian-like with separable means, we can interpret and manipulate the neurons in a more precise manner instead of zeroing out the neurons. Specifically, we only zero out a salient neuron (by saliency ranking) when its value is within a certain range. The core part of this design is selecting a suitable range that is highly correlated to the target concept $c$ that we want to remove. Intuitively, this range-based method conducts fine grained manipulation of neurons and thus alleviates the negative impact on other non-target concepts.
In this work, we adopt a straightforward choice of the aforementioned range, which is easy to calculate and implement. Specifically, we first calculate the empirical average $\mu\in\mathds{R}$ and standard deviation $\sigma\geq 0$ of the values of the salient neuron for all samples associated with the target concept $c\in C$. After that, we assign the range as $[\mu-\tau\times\sigma,\ \mu+\tau\times\sigma]$, where $\tau>0$ is a hyperparameter to make a tradeoff between erasing the target concept $c$ (using larger $\tau$) and smaller impact on other concepts (using smaller $\tau$). For this work, we use $\tau=2.5$. We chose this value to cover the majority of the target concept activations to produce a similar deterioration of the targeted concept.
Mathematically, our range-based method can be described as follows. To erase a target concept $c$ for a sample $x$ (e.g., a sentence or a word), we manipulate the salient neuron $h_{j}^{l}(x)$ ($j$-th element of the hidden state vector $\mathbf{h}^{l}(x)\in\mathds{R}^{d}$ at layer $l$).
$h_{j}^{l}(x)\begin{cases}=0&\text{ if }h_{j}^{l}(x)\in\text{CorrelatedRange}(l,j,c),\\ =h_{j}^{l}(x)&\text{ otherwise},\end{cases}$
where in this work we let
$\text{CorrelatedRange}(l,j,c)=[\mu-2.5\sigma,\ \mu+2.5\sigma],$
$\mu=\frac{1}{\left|H^{l}_{c}\right|}\sum_{\mathbf{h}^{l}\in H^{l}_{c}}h_{j}^{l},\ \quad\sigma=\sqrt{\frac{1}{\left|H^{l}_{c}\right|}\sum_{\mathbf{h}^{l}\in H^{l}_{c}}\left(h_{j}^{l}-\mu\right)^{2}}.$

### v2

@@ 4 Neuronal Activation Patterns
To quantitatively assess normality, while accounting for practical significance, we employ the KS test with an effect size threshold of 10%. This approach tests whether the distribution remains within a reasonable bound of normality, rather than testing for perfect normality, which is overly strict for real-world data. For each neuron, we normalize the activations to zero mean and unit variance, then compute the KS statistic against a standard normal distribution. The KS statistic represents the maximum absolute difference between the empirical and theoretical cumulative distribution functions. Using a threshold of 0.1 (allowing a maximum 10% deviation from normal), we find that close to 100% of the neurons exhibit practically normal distributions. The combination of near-ideal skewness and kurtosis values, visual confirmation through KDEs, and our effect size-based KS tests provide strong evidence that the activations follow approximately normal distributions.
We report quantitative statistics for all layers in Appendix Section I.1, which show that as layer depth increases, kurtosis steadily converges toward the Gaussian benchmark of 3.0, skewness remains near zero, and the 10% practical-normality score stays close to 1 across the network. A qualitative, layer-wise examination in Appendix Section I.2 further reveals that while all layers exhibit class level Gaussian-like activations. Early layers show substantial overlap between classes, this is consistent with the understanding that earlier layers focus on low-level features, not high level features like class. Beginning as early as layers 5–6, distinct class-specific Gaussians emerge and become progressively more separable in deeper layers, indicating a transition toward higher-level semantic representations.
5 Activation Ranges-guided Concept Erasure
Given that neuronal activations exhibit approximately Gaussian-like distributions with separable means, we can interpret and intervene on neurons more precisely than by ablating entire units. Specifically, $\mathsf{NeuronLens}$ ablates salient neurons identified through saliency ranking only when their activation falls within a selected range. The key idea is to identify a range that is strongly associated with the target concept $c$ intended for erasure. This range-based approach enables fine-grained ablation, thereby reducing unintended interference with non-target concepts. To validate our approach, we evaluate the causal efficacy of our method relative to neuron ablation using concept erasure experiments and assess the model’s latent capabilities following this intervention.
To calculate the aforementioned range, the framework utilises the means and standard deviations of the neuron activations. Specifically, first the empirical average $\mu\in\mathds{R}$ and standard deviation $\sigma\geq 0$ of the values of the salient neuron for all samples associated with the target concept $c\in C$ are calculated. After that, range is assigned as $[\mu-\tau\times\sigma,\ \mu+\tau\times\sigma]$, where $\tau>0$ is a hyperparameter to make a tradeoff between erasing the target concept $c$ (using larger $\tau$) and smaller impact on auxiliary concepts and general LLM capabilities (using smaller $\tau$). For this work, $\tau$ is set to $\tau=2.5$, assuming a fully Gaussian distribution. This threshold corresponds to a coverage of approximately $98.76\%$ of the distribution’s values, providing a slightly conservative bound for range-based interventions. Ablations for varying the hyperparameter $\tau$ are presented in Appendix Section H, the results indicate that targeted concept deteriorates up to 2.4-2.7 $\tau$ then plateaus, while auxiliary concepts begin to degrade further.
$h_{j}^{l}(x)=\begin{cases}\phi(x)&\text{if }h_{j}^{l}(x)\in\text{CR}(l,j,c)\\ h_{j}^{l}(x)&\text{otherwise}\end{cases}$
$\text{CR}(l,j,c)=[\mu\!-\!2.5\sigma,\ \mu\!+\!2.5\sigma],$
$\mu=\frac{1}{|H^{l}_{c}|}\sum_{\mathbf{h}^{l}\in H^{l}_{c}}h_{j}^{l},\sigma=\sqrt{\frac{1}{|H^{l}_{c}|}\sum_{\mathbf{h}^{l}\in H^{l}_{c}}(h_{j}^{l}-\mu)^{2}}$
where CR represents Correlated Range and $\phi()$ is the activation intervention function, which returns zero for the results presented in the main paper.
5.1 Experimental Setup
Models. This study employs both encoder and decoder-based models, including fine-tuned BERT (Devlin et al., 2019), DistilBERT (Sanh et al., 2020), GPT-2 (Radford et al., 2019), and pretrained Llama-3.2-3B (Grattafiori, 2024). We incorporate our methodology at the penultimate layer; ablation for layer selection is provided in the Appendix Section I. The training details for the models are provided in Appendix Section E.
For trained models (BERT, DistilBERT, and GPT-2), a higher proportion of neurons (up to 50%) can be ablated with a relatively minor impact on primary task performance and minimal interference with auxiliary concepts. This suggests substantial neuronal redundancy, wherein multiple neurons appear to encode overlapping features.
Datasets. We consider various classification based tasks; sentiment analysis (IMDB, (Maas et al., 2011)), (SST2, (Socher et al., 2013)), emotion detection (Dair-Ai/Emotions Saravia et al. (2018)), news classification (AG-News (Zhang et al., 2015)) and article content categorization (DBPedia-14 (Zhang et al., 2015)).
Table 3: Evaluation of selected models on IMDB, SST2, AG-News, and DBPedia-14 datasets using activation range and neuron masking techniques. Performance metrics are calculated using class level 10% trimmed means at the class level. Metrics are detailed in Section 2.1. For GPT-2 and Bert 50% and for Llama-3.2-3B 30% neurons selected.
@@ 5.1 Experimental Setup
5.2 Results and Analysis
Table 3 presents results for the concept removal task across five benchmark datasets (Class-wise detailed results are provided in Appendix Section J), demonstrating the effectiveness of our range-based masking approach compared to traditional neuron masking.
On binary classification tasks (IMDB, SST2), both masking approaches show moderate performance drops in targeted concepts. This suggests higher redundancy for coarser binary concepts. Multi-class classification tasks with fine-grained labels, such as AG-News, Emotions, and DBPedia-14, exhibit more pronounced effects under intervention. Range-based masking results in significant degradation of primary task performance while preserving auxiliary concept accuracy, this is particularly evident in results for AG-News.
GPT-2, despite being fine-tuned but trained in an autoregressive manner, shows substantially higher vulnerability with major drops in AG-NEWS ($\Delta_{acc}=-0.849$) and DBPedia-14 ($\Delta_{acc}=-0.782$). This increased sensitivity may be attributed to its autoregressive training objective, which potentially leads to more sequential and less redundant concept encodings. The Llama-3.2-3B model, evaluated in a few-shot setting without task-specific training, experiences the most severe degradation across all datasets (often exceeding $-0.90$), suggesting that pre-trained representations without task-specific fine-tuning are more vulnerable to targeted neuron interventions.
Table 4 in Appendix Section C presents the results showing the impact of concept erasure intervention on latent LLM capabilities such as fluency and generalization. Neuron masking degrades performance, increasing perplexity by (3.8-5.74) and lowering MMLU accuracy. In contrast, activation range masking raises perplexity by (0.5-1.1) points only, while preserving or improving MMLU scores indicates more precise and less disruptive removal.
Alternative activation interventions, beyond zeroing out, are explored in Appendix Section D, including the dampening method (Suau et al., 2024) and mean replacement (Suau et al., 2021). While these methods aim to manipulate without moving too far from the original representation, they exhibit limitations when applied to neurons. Specifically, neuron dampening increases perplexity by 2.9–3.7 points and often degrades MMLU accuracy (up to -0.045), whereas range-based dampening confines perplexity increases to 0.5–0.8 points and occasionally improves MMLU (up to +0.035). Similarly, mean replacement leads to substantial degradation when applied to neurons (perplexity increases of 7.4–8.8), while range-restricted replacement reduces the impact to below 0.7 points.
However, all approaches suffer from rigid static suppression or substitution, which fail to account for concept-specific activation dynamics. To address this issue, we introduce a novel adaptive dampening technique. This method modulates suppression in proportion to each activation’s deviation from its class-conditional mean, enabling data driven suppression. Adaptive dampening achieves the strongest balance across all metrics: perplexity remains low (0.41–0.61), MMLU is maintained or improved (up to +0.03), and collateral damage to auxiliary concepts is minimized (CAcc drops consistently below -0.3, often under -0.15), outperforming dampening, mean replacement and zeroing out approaches.
These results demonstrate that precise intervention in specific activation ranges, enables significantly more targeted concept manipulation while preserving auxiliary concepts, highlighting how conceptual information is encoded within specific activation patterns rather than isolated to individual neurons and underscoring the importance of activation ranges in capturing neuron-concept relationships.
Figure 6: Accuracy comparisons between Neuronal Range manipulation (green) and complete neuron manipulation (orange) methods on GPT-2 model.
Percentage Masking Effect
As more neurons are masked, performance gains of range-based masking over the neuron masking baseline become increasingly evident. Beyond a critical threshold of the number of masked neurons, baseline performance degrades sharply, while our method remains stable up to masking of 100% neurons. This arises from two factors: (1) models have a large number of polysemantic neurons and higher masking rates increase the chance of ablating them, and (2) blocking/manipulating a higher percentage of the model’s neurons creates a significant deviation from the original model’s behavior. For low-activation neurons with respect to the concept of interest, discrete neuron masking i.e. completely masking out a neuron, becomes unreliable, as shown in Figure 6, with a steep performance drop after masking 50% neurons. This underscores the need for finer-grained attribution; our range-based method offers such precision, preserving model behaviour under extensive masking.
The relatively stable results on auxiliary concept when using range-based masking at high percentage of neurons reduces the need to find an optimum threshold for the number of neurons to ablate which is critical to neuron masking.
6 Related Work
While we have discussed closely related approaches in Section 2, here, we briefly review additional relevant techniques. Circuit discovery identifies groups of neurons that jointly encode concepts, providing a structured view of model behavior (Marks et al., 2024; Conmy et al., 2023; Olah et al., 2020). However, extracting circuits is computationally intensive and lacks fine-grained neuron-level attribution. Gradient-based methods attribute predictions to input features by tracking gradients through the network, with integrated gradients (Sundararajan et al., 2017; Dai et al., 2022) being a widely used approach. However, they struggle with polysemanticity, as they do not disentangle overlapping concepts within neurons. Causal analysis methods intervene on internal components to assess their role in encoding concepts. Causal tracing measures the effect of corrupting activations on model performance (Vig et al., 2020; Meng et al., 2022), while causal mediation analysis quantifies information propagation through neurons (Vig et al., 2020). Although effective, these techniques require costly perturbation experiments. Beyond neuron-level analysis, representation-level methods examine hidden states and their relationship to model outputs and concepts (Veldhoen et al., 2016; Tenney et al., 2019; Liu et al., 2019). Sparse probing (Gurnee et al., 2023) compresses hidden representations into sparse, interpretable subspaces. While prior work has advanced interpretability, most methods rely on discrete neuron-to-concept mappings, which fail to account for polysemanticity (Sajjad et al., 2022). Our work extends activation-based approaches by introducing activation ranges as the unit of interpretability to enable more precise concept attribution and intervention.
7 Conclusion
In this work, we challenged traditional assumptions about neuron interpretability by reframing polysemanticity as a resource rather than a limitation in interpreting neurons. Through an in depth analysis, we uncovered that neuronal activations for individual concepts exhibit distinct, Gaussian-like distributions. This discovery allows for a more precise understanding of how neurons encode multiple concepts, enabling us to move beyond coarse, monolithic neuron-to-concept mappings. Building upon these insights, we proposed $\mathsf{NeuronLens}$, a novel range-based framework for neuron interpretation and manipulation. $\mathsf{NeuronLens}$ offers fine-grained control that reduces interference with unrelated concepts by attributing specific activation ranges within neurons to individual concepts. Extensive empirical evaluations demonstrated that $\mathsf{NeuronLens}$ outperforms neuronal attribution methods in maintaining concept-specific precision while minimizing unintended side effects. Notably, while targeted concept removal remains equally effective when comparing neuron vs range based interventions, our approach achieves superior preservation of auxiliary concepts without compromising the primary goal. An important direction for future work is exploring our range-based method as a metric for quantifying polysemanticity in neural networks. This approach may also serve as a diagnostic tool to evaluate the effectiveness of sparse autoencoders (SAEs) in disentangling concept representations across individual neurons.
Appendix A Impact Statement
Appendix B Limitations
While $\mathsf{NeuronLens}$ can disentangle polysemanticity to a degree using Gaussian Like Distribution, it is unable to completely disentangle concepts encoded in the polysemantic neurons, because there still is a significant overlap in the distributions of concepts in activations. Additionally, in this work, we use $\tau$ to be a fixed value of 2.5 to make the comparison of approaches fair, but $\tau$ selection can be optimized to be more sophisticated. We also get results primarily from the penultimate layer, and not the intermediate or earlier layers, however, we do give ablation and rationale for this choice in Appendix I
Appendix C General LLM Capabilities
We evaluate the general capabilities of large language models (LLMs) using the MMLU benchmark [Hendrycks et al., 2021] and perplexity scores on Wikipedia texts [Foundation, ]. Table 4 presents the comparative performance of neuron masking and activation range masking. Neuron masking leads to notable increases in perplexity, exceeding 3 points in the best case, whereas range masking results in a maximum increase of only 1.1. In terms of MMLU accuracy, neuron masking consistently reduces performance across all settings, while range masking preserves or improves performance in most cases, with degradation observed in only one instance.
Table 4: Evaluation of LLMs latent capabilities using Wikipedia for perplexity and zeroshot MMLU for testing generalisation capabilities. Llama-3.2-3B model
Dataset Base Values Neuron Masking Activation Range Masking
Perplexity MMLU Perplexity MMLU Perplexity MMLU
IMDB 7.007 0.530 10.990 0.515 7.550 0.530
SST2 7.007 0.530 11.688 0.510 8.150 0.537
AG-NEWS 7.007 0.530 12.757 0.510 8.022 0.533
Emotions 7.007 0.530 11.630 0.526 8.063 0.526
DBPedia-14 7.007 0.530 12.230 0.507 7.903 0.535
Appendix D Activation Intervention
In the main text, we primarily presented results using a “zeroing out” strategy for neuron manipulation. This approach was chosen to compare neuron manipulation against range-based manipulation. However, zeroing out is considered a suboptimal strategy [Suau et al., 2024]. The primary concern with standard zeroing-out approaches is that they distort the activation distribution significantly, diverging from that of the original model. However, our range-based method selectively zeroes out only a narrow slice of the activation spectrum, thereby mitigating the adverse effects associated with hard erasure.
In this section, we explore alternative, more optimized strategies for concept removal. We also introduce a novel range-based scaling strategy that has demonstrated superior results.
Below, we explore various activation intervention strategies, comparing traditional neuron-level approaches with the nuanced range-based technique. Our comprehensive evaluation reveals that range-based manipulations consistently outperform neuron interventions across multiple metrics, with significantly less disruption to the model’s general capabilities.
Among all techniques examined, our novel adaptive dampening approach emerges as the most effective, maintaining targeted concept suppression while minimizing collateral impact on auxiliary concepts and preserving overall language modelling capabilities. This pattern holds true across different intervention methods including zeroing out, dampening, and mean replacement strategies.
D.1 Dampening
In their work, Suau et al. [2024] propose using a dampening function rather than setting neuron activations to zero outright. This approach, referred to as DAMP, corresponds to a specific choice of the intervention function $\phi(x)=\alpha x$, where $0\leq\alpha\leq 1$. In this formulation, the activations of selected neurons are scaled down by a factor $\alpha$ instead of being completely suppressed. Here, $x$ represents neuron activation. The rationale behind dampening is that a fixed intervention (like zeroing out) can disrupt the LLM’s inference dynamics, especially when a large number of neurons ($k$) are involved, thereby limiting its effectiveness. Dampening offers a less destructive intervention by allowing contextual signals to continue passing through the network. This, in turn, permits intervention on a larger set of expert neurons, potentially achieving stronger mitigation of the targeted concept.